2. x + 6x + 18x sin 2t, x(0) = -1, x0) = 1.

Use the Laplace transform to solve initial value problems
Answer #1

x(0)=1 -) x(6)=-1, x +6xt x +18% Sin QE) domain converting into Laplace a sex(3) - 5X(0) - x(0) + 6/8 x(3) -R0)] +18 x(383 2 [A+c] + 8? [GA+B+0] +8 [18 A+ +63+4c]t 8 18 BtUD Atc=0 GAI BID=0 18 A +68+ UC=0 - A=-c 269)+82-0 -18A+6B - ua=0 8 IUA +62(3) ㅗ Sin (20) + Bécs (36) + 336 38+3 (853) +9 (st3y+9 I.LT ) ☺-@ -3 Col (24) + ? Sin (3.6). 85 85 est cos(3E) + 2/3 Sin (3t